OpenStudy (anonymous):

A solution of salt and water contains 50 grams of water per 100 milliliters of the solution. If 1 mole of water weighs 16 grams, _____ moles of water would be present in 25 milliliters of the solution. [Round off your answer to two decimal places]

OpenStudy (shinalcantara):

use ratio and proportion

OpenStudy (anonymous):

I honestly don't know the answer to this and I'm trying to finish two module tests before I get withdrawn.. :(

OpenStudy (anonymous):

And I need help with a few more :/

OpenStudy (shinalcantara):

\[\frac{ 50 }{ 100 } = \frac{ x }{ 25 }\]

OpenStudy (shinalcantara):

now solve for x. x is the number of grams of water

OpenStudy (anonymous):

12.5?

OpenStudy (shinalcantara):

\[\frac{ 1 }{ 16 } = \frac{ y }{ 12.5 }\]

OpenStudy (anonymous):

I'm lost now

OpenStudy (shinalcantara):

since you have 12.5 as your answer, you try substitute it for another ratio & proportion that will indicate the number of moles..

OpenStudy (shinalcantara):

try to reread the problem and understand the ratio & proportion i'm talking about

OpenStudy (anonymous):

I'm so confused.

OpenStudy (anonymous):

I'm sorry, algebra is my worst subject.

OpenStudy (cj49):

50gms of water per 100 mil so how much grm of water wud 1 mil contain

OpenStudy (anonymous):

0.78125?

OpenStudy (shinalcantara):

in finding the number of grams of water for a 25mL solution, we'll refer it to the given data that 50 grams does have 100 mL using ratio & proportion

OpenStudy (cj49):

1 gram water per 2 ml solution

OpenStudy (cj49):

50/100=0.5(per mil)

OpenStudy (shinalcantara):

yep but you're asked to have it rounded to two decimal places

OpenStudy (cj49):

12.5 grams water per 25 ml solution

OpenStudy (cj49):

1mole= 16gms

OpenStudy (cj49):

1 16 x 12.5

OpenStudy (cj49):

12.5/16 =?

OpenStudy (anonymous):

That equals 0.78125 right?

OpenStudy (anonymous):

I don't get the two decimal part..

OpenStudy (cj49):

the ans wud be 0.78

OpenStudy (shinalcantara):

the answer would be 0.78

OpenStudy (anonymous):

Can you help with a few more?

OpenStudy (shinalcantara):

the two decimal places as it is asked in the problem indicates that you're to round it off such that there are only two numbers after the decimal point

OpenStudy (shinalcantara):

or to the right of the decimal point

OpenStudy (anonymous):

Okay! The regular price of a child's entry ticket to a water park is $6 less than that for an adult's. The park offers half off all entry tickets during the off-peak season. The Jordans paid a total of $69 for 2 adult tickets and 1 child's ticket to the water park during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket. : 69 = x + one-half(x - 6) What is the regular price of a child's ticket? $38 $42 $48 $66

OpenStudy (cj49):

the ans is 42

OpenStudy (anonymous):

Thank you! The radius of a cone-shaped tank is 4 feet less than its height. If the height of the tank is (x - 3) feet, the expression below shows the volume of the tank: 1 over 3π (x - 7)2 (x - 3) What does the factor π(x - 7)2 (x - 3) represent? The area of the base of the tank The area of the curved sides of the tank The volume of 3 of the same cone-shaped tanks The volume of 6 of the same cone-shaped tanks

OpenStudy (anonymous):

the question mark things are the pi sign

OpenStudy (cj49):

sorry mate geometry is not my thing

OpenStudy (anonymous):

Simplify the following expression x - 4y + 5z - 3x - y + 2z = ?. 2x - 5y + 7z -2x - 5y - 7z -2x - 5y + 7z -2x -5y + 3z

OpenStudy (cj49):

-2x-5y+7z

OpenStudy (anonymous):

Yay got it right haha. Evaluate: -3 - (-6). -3 -18 -9 3

OpenStudy (cj49):

-3+6=3

OpenStudy (anonymous):

Which table best represents the steps to solve the equation x over 8 - 5 = -1? Step Justification Step 1: x over 8 = -1 + 5 The same number + 5 can be added to both sides of the equation Step 2: x over 8 = 4 Adding -1 and 5 on the right hand side Step 3: x = 4 • 8 The same number 8 can be multiplied on both sides of the equation Step 4: x = 32 Multiplying 4 and 8 on the right hand side Step Justification Step 1: x over 8 = -1 - 5 The same number - 5 can be added to both sides of the equation Step 2: x over 8 = -6 Adding -1 and -5 on the right hand side Step 3: x = -6 • 8 The same number 8 can be multiplied on both sides of the equation Step 4: x = -48 Multiplying -6 and 8 on the right hand side Step Justification Step 1: x over 8 = -1 + 5 The same number - 5 can be added to both sides of the equation Step 2: x over 8 = 4 Adding -1 and 5 on the right hand side Step 3: x = 4 - 8 The same number 8 can be subtracted on both sides of the equation Step 4: x = -4 Subtracting 8 from 4 on the right hand side Step Justification Step 1: x over 8 = -1 + 5 The same number + 5 can be added to both sides of the equation Step 2: x over 8 = 4 Adding -1 and 5 on the right hand side Step 3: x = 8 ÷ 4 The same number 8 can be divided on both sides of the equation Step 4: x = 2 Subtracting 8 from 4 on the right hand side

OpenStudy (shinalcantara):

the volume of 3 of the same cone-shaped tanks

OpenStudy (cj49):

u mean x/8 -5=-1

OpenStudy (cj49):

the 1st one is the correct method

OpenStudy (anonymous):

Okay, thank you so much!!

OpenStudy (cj49):

np mate c:

OpenStudy (anonymous):

Rita's company reimburses her expenses on food, lodging, and conveyance during business trips. The company pays $55 a day for food and lodging and $0.45 for each mile traveled. Rita drove 300 miles and was reimbursed $2,335. Part A: Create an equation that will determine the number of days x on the trip (3 points) Part B: Solve this equation justifying each step with an algebraic property of equality. (6 points) Part C: How many days did Rita spend on this trip? (1 points)